LABORATORY DETECTION OF THE TRANS-GAUCHE CONFORMER OF ETHYL FORMATE.

\underline{JUSTIN L. NEILL}, MATT T. MUCKLE, DANIEL P. ZALESKI, AND BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., P.O. Box 400319, Charlottesville, VA 22904; V. LATTANZI, S. SPEZZANO, AND M.C. MCCARTHY, Harvard-Smithsonian Center for Astrophysics, 60 Garden St., Cambridge, MA 02138, and School of Engineering & Applied Sciences, Harvard University, 29 Oxford St., Cambridge MA 02138.

Ethyl formate has two coordinates of conformational flexibility, in the ester (O=C-O-C) and ethyl (C-O-C-C) dihedral angles. Two conformers, one with a cis ester and trans ethyl orientation, the other with a cis ester and gauche ethyl orientation, have been previously detected by rotational spectroscopy.^{a} In addition, the cis-trans isomer, the lowest-energy conformer, has recently been detected in the SgrB2(N) hot core.^{b} The third conformer of ethyl formate, with a trans ester orientation and gauche ethyl orientation, is significantly higher in energy than the cis-trans conformer according to electronic structure calculations (1900 cm^{−1}, or 1330 K), but there is a barrier of 2870 cm^{−1} (2000 K) for this conformer to relax into the more stable cis ester potential well, and so local thermodynamic equilibrium between these conformers is not expected in the interstellar medium. Similar behavior is found for the trans ester conformer of methyl formate, for which a tentative detection in SgrB2(N) was presented at this meeting last year, with a column density roughly 1% of that of the more stable cis isomer.^{c} Here we report the laboratory detection of trans-gauche ethyl formate using Fourier transform microwave spectroscopy; its low population at room temperature equilibrium has been enhanced by the use of a pulsed discharge nozzle. The spectrum is complicated by a low barrier (140 cm^{−1}) to tunneling between equivalent structures.

aJ.M. Riveros and E.B. Wilson, J. Chem. Phys. 46, 4605 (1967).

bA. Belloche. et al., A&A 499, 215 (2009).

cM.T. Muckle et al., RH15, 64th International Symposium on Molecular Spectroscopy (2009).